# Assessing the star formation history of all-sky and part-sky 100 pc white dwarf samples

### v1

@@ 2.1.2 Gaia Catalogue of Nearby Stars sample
The GCNS is a Gaia Early Data Release 3 catalogue laid out in Gaia Collaboration et al. (2021). They define the Solar neighbourhood to be all objects within a sphere of radius 100 pc centred on the Sun, and select objects via a random forest classification algorithm using various parameters defined in Gaia such as the renormalised unit weight error and the parallax error. The selected sample is roughly uniform over the sky, as expected, and most rejected objects are in regions of high surface density, also as expected. To extract the white dwarf only sample, another white dwarf random forest classifier was created using the three Gaia photometric bands, proper motions, parallaxes and parallax uncertainties. The full catalogue contains over 330 000 stars, and over 15 000 white dwarfs within 100 pc of the Sun.
@@ 2.1.3 GSPC-WD sample
A new feature of Gaia DR3 compared to DR2 or EDR3, is the inclusion of flux-calibrated low resolution spectrophotometry for over 200 million sources (Gaia Collaboration et al., 2023). The raw spectra in the blue (BP) and red (RP) photometric channels are calibrated, then merged into a single XP spectrum (Montegriffo, P. et al., 2023; De Angeli, F. et al., 2023). Vincent et al. (2024) created an apparent magnitude limited (G<20.5<20.5) catalogue of white dwarfs (the Gaia Synthetic Photometry Catalogue for White Dwarfs, GSPC-WD) with XP spectra and generated SDSS-like photometry for each white dwarf. A machine learning based classification system gave each spectrum a spectral type and composition, from which white dwarf parameters could be calculated using the same data and tailored model atmospheres. The catalogue’s completeness follows that of Gentile Fusillo et al. (2021) up to about 50 pc, beyond which the completeness of the GSPC-WD drops below that of Gentile Fusillo’s catalogue, with an effective magnitude limit of about $G\approx 20$. There is also evidence of a colour-dependent decrease in completeness towards redder colours in the GSPC-WD (Jiménez-Esteban et al., 2023). By using this catalogue, the number of white dwarfs with known spectral type increases by a factor of about 3 within 100 pc (Vincent et al., 2024).
By keeping only objects with a probability of being a white dwarf greater than 0.75, limiting the sample to only those within 100 pc and limiting objects to those with a white dwarf mass between 0.541.340.54-1.34 M to eliminate binary white dwarfs, the GSPC-WD contains 13 195 white dwarfs with spectral classifications, masses, $T_{\rm eff}$, $\log g$, and more, which we use directly from the Vincent et al. (2024) source catalogue. We maintain the white dwarf probability cut despite these objects having white dwarf spectral types to keep the selection criteria the same across all samples. This removes less than 1 per cent of their catalogue.
@@ 2.2 Part-Sky 100 pc Samples
Figure 3: Top: The footprint in RA and Dec (degrees) of the SDSS sample taken from Kilic et al. (2025), and the simplified footprint used to generate our simulation of the SDSS sample. Our intention was not to match the footprints completely, but mostly simulate the same parts of the sky. Middle: The footprint in RA and Dec of the DESI sample taken from DESI Collaboration et al. (2025). Bottom: The simulated 4MOST footprint based on the planned sky coverage of the instrument.
Not all surveys cover and observe the whole of the sky; for instance ground-based MOS surveys are restricted to either the northern or southern hemispheres based on the telescope location, and often have limited footprints, for instance by focusing (or avoiding) the Galactic plane regions as driven by other Galactic (or extra-galactic) science cases. Furthermore, ground MOS surveys typically have a lower completeness than photometric surveys, owing to both the instrumental limitations in crowded fields, and the pressure factor of different science sub-surveys trying to obtain limited fibre time.
@@
Here, we assess the white dwarf spectroscopic samples of Kilic et al. (2025) based on the SDSS I-IV footprint, the first DESI data release (DESI Collaboration et al., 2025), and we predict the footprint and white dwarf sample that is expected to be available after the completion of the 4MOST 100PC and white dwarf sub-surveys around 2031 (de Jong et al., 2019).
@@ 2.2.1 SDSS
Kilic et al. (2025) use SDSS DR9 and Pan-STARRS DR1 photometry matched with Gaia DR3 to create a catalogue of white dwarfs within 100 pc. They focussed on reliability rather than completeness in their sample, and on spectroscopic follow up. By combining previous literature spectral classifications and follow up observations conducted in their work, the sample has 75 per cent spectroscopic completeness for white dwarfs identified in the SDSS DR9 footprint. Their completeness reaches 86 per cent for $T_{\rm eff}>5000$ K, illustrating a bias against cooler, redder white dwarfs similarly to the GSPC-WD sample.
Of the 3145 objects in the sample, we keep 2744 once we have selected only objects with a probability of being a white dwarf greater than 0.75, and only those with a white dwarf mass between 0.541.340.54-1.34 M. The probability cut is kept despite objects in this sample being confirmed white dwarfs to keep the selection criteria the same for all samples used. The $P_{\rm WD}$ used for each object is that of Gentile Fusillo et al. (2021). The sky footprint of this sample can be seen in the top left of Fig. 3 compared with the approximated analytical footprint used in our population synthesis simulations in the top right.
@@ 2.2.2 DESI
The Dark Energy Spectroscopic Instrument (DESI) Data Release 1 (DESI Collaboration et al., 2025) provides spectra for around 18 million unique objects and is mainly planned to observe distant galaxies and quasars to explore the nature of dark energy. However, during periods affected by lunar light contamination or twilight, the survey will target nearby bright galaxies and stars, including white dwarfs, as part of its ‘bright’ programme. It employs a less biased spectroscopic target selection (Manser et al., 2024) than the SDSS I–IV catalogues. Based on the input catalogue of Gentile Fusillo et al. (2019), DESI chooses white dwarf targets largely at random from its input catalogue, within its field of view and magnitude limits, immediately prior to each observation. The full DESI DR1 white dwarf sample contains over 41 000 white dwarfs at high spectroscopic signal-to-noise and resolution of $R\approx 2000-5000$. When we limit the sample to those within 100 pc, with a probability of being a white dwarf above 0.75, and within the mass range 0.541.340.54-1.34 M, the DESI DR1 100 pc sample contains 2978 white dwarfs. The sky footprint of this sample can be seen in the middle left of Fig. 3, compared with our simulated footprint in the middle right.
@@
The 4-metre Multi-Object Spectroscopic Telescope (4MOST) is a project by the European Southern Observatory to provide the VISTA telescope with a fibre-fed spectroscopic survey facility, with first light in October 2025. The aim is to provide complementary data for large scale sky surveys such as Gaia, Euclid, and PLATO. With the knowledge that 4MOST plans to survey most of the southern sky, we can begin to predict the data to be obtained. 4MOST will survey all RA values, but will be limited to between 5 to -70 degrees in Dec. The expected and simulated footprint of the 4MOST sample can be seen at the bottom of Fig. 3.
The 4MOST 100PC sub-survey (PI Tremblay), within the S3 - Milky Way Bulge and Disk Low Resolution Survey (4MIDABLE-LR) survey (Chiappini et al., 2019), is dedicated to obtaining optical medium-resolution ($R\approx$ 5000) spectra of $\approx$200 000 Gaia-selected main-sequence stars and $\approx$8000 white dwarfs within the 4MOST footprint and 100 pc of the Sun. One goal of the survey is to compare the SFH derived from stars and white dwarfs, and calibrate stellar ages. There is also a 4MOST white dwarf sub-survey (PIs Tremblay/Gentile Fusillo) set to observe white dwarfs at all distances. Both sub-surveys have a magnitude limit of $G<19.5$, and aim for a signal-to-noise of 30 at 4500 Å.
@@ 3.1 Simulation
We use the single-star population synthesis code of Roberts et al. (2025) to generate populations of stars in the local Galaxy according to the initial mass function, a given SFH, and a spatial distribution within the disc. They are then evolved kinematically and temporally, where it then selects the current population of white dwarfs remaining within 100 pc at the present day. The code generates the magnitudes, luminosities, and other observational properties of all the white dwarfs in the simulation to compare against the observed distributions of the samples discussed in the previous section. The simulation code takes an assumed SFH form. We use the same four assumed approximate forms as Roberts et al. (2025), shown in Fig. 7 of that paper.
Each ingredient in the simulation introduces an uncertainty into the final predicted ages, luminosities, or magnitudes of the white dwarfs. The ingredients and their uncertainties are listed in Table 2, with both the original size of the uncertainties and the reduced uncertainties after fitting simultaneously the white dwarf mass distribution and local SFH in Roberts et al. (2025). In this work, all uncertainties use the reduced sizes. Some of these uncertainties are systematic (population age, initial mass function, main sequence lifetimes, spectral evolution, merger delays, age vs kinematic relation, cooling ages, crystallisation delays) and some affect individual stars in each simulation (metallicity, and any Gaia data errors associated with each star). We refer the reader to Roberts et al. (2025) for further details.
@@
This work focusses on two of the three methods used in Roberts et al. (2025), comparisons of the white dwarf luminosity function and absolute Gaia G magnitude distribution, with the third - the direct age method - explored only in Appendix A for those who wish to see all three methods tested on the 100 pc samples.
@@ 3.2 Luminosity function
To calculate the bolometric magnitudes of white dwarfs in the sample, the effective temperature and radius of the white dwarf are required. This itself requires knowing the atmospheric composition of the white dwarf, i.e. hydrogen- or helium-dominated, as fitting temperature and radius requires fitting a model atmosphere and mass-radius relation to the photometry and astrometry (or, although not used in this case, fitting the same models to spectroscopy). With the 40 pc sample, all of this composition information is known, but for the 100 pc samples used in this work, this is only known for the samples of Vincent et al. (2024) and Kilic et al. (2025). Because of this, we chose to use temperatures and masses as derived from pure-hydrogen atmospheres for samples where no spectroscopic or incomplete spectroscopic data is known. Fig. 2 demonstrates that this assumption makes little difference, as the derived 100 pc white dwarf luminosity functions are very similar to the one for the volume complete 40 pc sample, which has full atmospheric composition information. We note that most cool white dwarfs ($T_{\rm eff}<5000\,$K) are of DC spectral type with unconstrained compositions (e.g. Elms et al., 2022), hence the assumption of pure-hydrogen atmospheres is actually done for all samples at the faint end of the luminosity function, regardless of the availability of spectroscopy.
On the other hand, in the simulated luminosity functions, all samples without composition information rely on the 25 per cent helium fraction as described in Table 2 and each white dwarf is randomly assigned a composition. For the samples of Vincent et al. (2024) and Kilic et al. (2025), we can predict a slightly more accurate luminosity function using the known compositions of each white dwarf. We find that these two options make very little difference in the predicted luminosity functions.
Figure 4: Top: Results from the luminosity function method with population synthesis using all four assumed star formation forms (constant; Roberts et al. 2025 direct age method; Fantin et al. 2019; Mor et al. 2019) compared to observed 100 pc white dwarf sample of Gentile Fusillo et al. (2021). Error bars are the standard deviation of 100 runs of the simulations, and Poisson errors for the observed sample. Bottom: Results from the absolute Gaia G magnitude method with all four assumed star formation forms compared to the observed 100 pc sample of Gentile Fusillo et al. (2021). Error bars follow the same convention as above.
@@ 3.3 Absolute Gaia G magnitude distribution
All observed magnitudes are dereddened based on either the $V$-band extinction ($A_{V}$) for that object in Gentile Fusillo et al. (2021), or if missing, an $A_{V}$ extinction value sampled from a Gaussian distribution with the mean as 0.02 mag (and standard deviation of 0.015 mag), which is the average extinction over the whole 100 pc sample (Gentile Fusillo et al., 2021; Sahu et al., 2024).
Table 2: Literature and reduced uncertainties derived in Roberts et al. (2025) as input of single-star population synthesis. The reduced uncertainties were used for all simulations in this work.
@@
Population age $\mu=10.6$ Gyr, $\sigma=0.7$ Gyr Cukanovaite et al. (2023) $\sigma=0.5$ Gyr
Initial mass function $\mu=2.35$, $\sigma=0.1$ Salpeter (1955) $\sigma=0.075$
Initial metallicity $\mu=0.0134$, $\sigma=0.0104$ Asplund et al. (2009) no change
Main sequence lifetimes multiplicative, $\mu=1$, $\sigma=0.048$ Byrne et al. (2025) $\sigma=0.035$
He-atmosphere WD fraction $\mu=0.25$, $\sigma=0.1$ O’Brien et al. (2024) $\sigma=0.065$
Merger delays off/on, equal weighting Roberts et al. (2025) off/on, 43/57 weighting
Age vs. kinematic relation multiplicative to slope of scale height, $\mu=1$, $\sigma=0.3$ Roberts et al. (2025) no change
Cooling models multiplicative, $\mu=1$, $\sigma=0.06$ Bédard et al. (2020) $\sigma=0.045$
Crystallisation cooling delay off/on, equal weighting Roberts et al. (2025) no change

### v2

@@ 2.1.2 Gaia Catalogue of Nearby Stars sample
The GCNS is a Gaia Early Data Release 3 catalogue laid out in GCNS_2021. They define the Solar neighbourhood to be all objects within a sphere of radius 100 pc centred on the Sun, and select objects via a random forest classification algorithm using various parameters defined in Gaia such as the renormalised unit weight error and the parallax error. The selected sample is roughly uniform over the sky, as expected, and most rejected objects are in regions of high surface density, also as expected. To extract the white dwarf only sample, another white dwarf random forest classifier was created using the three Gaia photometric bands, proper motions, parallaxes and parallax uncertainties. The full catalogue contains over 330 000 stars, and over 15 000 white dwarfs within 100 pc of the Sun.
@@ 2.1.3 GSPC-WD sample
A new feature of Gaia DR3 compared to DR2 or EDR3, is the inclusion of flux-calibrated low resolution spectrophotometry for over 200 million sources (GaiaDR3_2023). The raw spectra in the blue (BP) and red (RP) photometric channels are calibrated, then merged into a single XP spectrum (Montegriffo_2023; DeAngeli_2023). Vincent_2024 created an apparent magnitude limited (G<20.5<20.5) catalogue of white dwarfs (the Gaia Synthetic Photometry Catalogue for White Dwarfs, GSPC-WD) with XP spectra and generated SDSS-like photometry for each white dwarf. A machine learning based classification system gave each spectrum a spectral type and composition, from which white dwarf parameters could be calculated using the same data and tailored model atmospheres. The catalogue’s completeness follows that of GentileFusillo_2021 up to about 50 pc, beyond which the completeness of the GSPC-WD drops below that of Gentile Fusillo’s catalogue, with an effective magnitude limit of about $G\approx 20$. There is also evidence of a colour-dependent decrease in completeness towards redder colours in the GSPC-WD (JimenezEsteban_2023). By using this catalogue, the number of white dwarfs with known spectral type increases by a factor of about 3 within 100 pc (Vincent_2024).
By keeping only objects with a probability of being a white dwarf greater than 0.75, limiting the sample to only those within 100 pc and limiting objects to those with a white dwarf mass between 0.541.340.54-1.34 M to reduce the impact of binary white dwarfs, the GSPC-WD contains 13 195 white dwarfs with spectral classifications, masses, $T_{\rm eff}$, $\log g$, and more, which we use directly from the Vincent_2024 source catalogue. We maintain the white dwarf probability cut despite these objects having white dwarf spectral types to keep the selection criteria the same across all samples. This removes less than 1 per cent of their catalogue.
@@ 2.2 Part-Sky 100 pc Samples
Figure 3: Top: The footprint in RA and Dec (degrees) of the SDSS sample taken from Kilic_2025_100pc, and the simplified footprint used to generate our simulation of the SDSS sample. Our intention was not to match the footprints completely, but mostly simulate the same parts of the sky. Middle: The footprint in RA and Dec of the DESI sample taken from DESI_2025. Bottom: The simulated 4MOST footprint based on the planned sky coverage of the instrument.
Not all surveys cover and observe the whole of the sky; for instance ground-based MOS surveys are restricted based on the telescope location, and often have limited footprints, for instance by focusing (or avoiding) the Galactic plane regions as driven by other Galactic (or extra-galactic) science cases. Furthermore, ground MOS surveys typically have a lower completeness than photometric surveys, owing to both the instrumental limitations in crowded fields, and the pressure factor of different science sub-surveys trying to obtain limited fibre time.
@@
Here, we assess the white dwarf spectroscopic samples of Kilic_2025_100pc based on the SDSS I-IV footprint, the first DESI data release (DESI_2025), and we predict the footprint and white dwarf sample that is expected to be available after the completion of the 4MOST 100PC and white dwarf sub-surveys around 2031 (4MOST).
@@ 2.2.1 SDSS
Kilic_2025_100pc use SDSS DR9 and Pan-STARRS DR1 photometry matched with Gaia DR3 to create a catalogue of white dwarfs within 100 pc. They focussed on reliability rather than completeness in their sample, and on spectroscopic follow up. By combining previous literature spectral classifications and follow up observations conducted in their work, the sample has 75 per cent spectroscopic completeness for white dwarfs identified in the SDSS DR9 footprint. Their completeness reaches 86 per cent for $T_{\rm eff}>5000$ K, illustrating a bias against cooler, redder white dwarfs similarly to the GSPC-WD sample.
Of the 3145 objects in the sample, we keep 2744 once we have selected only objects with a probability of being a white dwarf greater than 0.75, and only those with a white dwarf mass between 0.541.340.54-1.34 M. The probability cut is kept despite objects in this sample being confirmed white dwarfs to keep the selection criteria the same for all samples used. The $P_{\rm WD}$ used for each object is that of GentileFusillo_2021. The sky footprint of this sample can be seen in the top left of Fig. 3 compared with the approximated analytical footprint used in our population synthesis simulations in the top right.
@@ 2.2.2 DESI
The Dark Energy Spectroscopic Instrument (DESI) Data Release 1 (DESI_2025) provides medium-resolution spectra ($R\approx$ 2000–5000) for around 18 million unique objects and is mainly planned to observe distant galaxies and quasars to explore the nature of dark energy. However, during periods affected by lunar light contamination or twilight, the survey will target nearby bright galaxies and stars, including white dwarfs, as part of its ‘bright’ programme (Cooper_2023). Using an analogue to the Gaia DR2 catalogue of GentileFusillo_2019 as input catalogue, DESI chooses white dwarf targets largely at random, within its field of view and magnitude limits, immediately prior to each observation. This results in a less biased spectroscopic target selection (Manser_2024) than the SDSS I–IV catalogues.
To define the DESI DR1 100 pc white dwarf sample, we downloaded the full DESI DR1 dataset [^1] https://data.desi.lbl.gov/public/dr1/ and performed a cross-match with the Gaia EDR3 catalogue of GentileFusillo_2021 using the unique Gaia DR2 identifier. This yielded a total of 43 440 unique sources with at least one DESI spectrum. Keeping only objects within 100 pc, with a probability $P_{\rm WD}\geq 0.75$, and within the mass range 0.541.340.54-1.34 M, we obtained a final sample of 2978 white dwarfs. The sky footprint of this sample can be seen in the middle left of Fig. 3, compared with our simulated footprint in the middle right.
@@
The 4-metre Multi-Object Spectroscopic Telescope (4MOST) is a project by the European Southern Observatory to provide the VISTA telescope with a fibre-fed spectroscopic survey facility, with first light in October 2025. The aim is to provide complementary data for large scale sky surveys such as Gaia, Euclid, and PLATO. With the knowledge that 4MOST plans to survey most of the southern sky, we can begin to predict the data to be obtained. 4MOST will survey all RA values, but will be limited to between 5 to -80 degrees in Dec. The expected and simulated footprint of the 4MOST sample can be seen at the bottom of Fig. 3.
The 4MOST 100PC sub-survey (PI Tremblay), within the S3 - Milky Way Bulge and Disk Low Resolution Survey (4MIDABLE-LR) survey (4MIDABLE), is dedicated to obtaining optical medium-resolution ($R\approx$ 5000) spectra of $\approx$200 000 Gaia-selected main-sequence stars and $\approx$8000 white dwarfs within the 4MOST footprint and 100 pc of the Sun. One goal of the survey is to compare the SFH derived from stars and white dwarfs, and calibrate stellar ages. There is also a 4MOST white dwarf sub-survey (PIs Tremblay/Gentile Fusillo) set to observe white dwarfs at all distances. The aim is for a signal-to-noise (S/N) of 30 at 4500 Å for $G<18.5$ mag, and a S/N of 15 at $18.5<G<19.5$. 4MOST will observe even fainter targets from its ancillary catalogue, but we use an effective $G=19.5$ mag cut-off corresponding to high-quality (high S/N) data.
@@ 3.1 Simulation
We use the single-star population synthesis code of Roberts_2025 to generate populations of stars in the local Galaxy according to the initial mass function, a given SFH, and a spatial distribution within the disc. They are then evolved kinematically and temporally, where it then selects the current population of white dwarfs remaining within 100 pc at the present day. The code generates the magnitudes, luminosities, and other observational properties of all the white dwarfs in the simulation to compare against the observed distributions of the samples discussed in the previous section. The simulation code takes an assumed SFH form. We use the same four assumed approximate forms as Roberts_2025, shown in Fig. 7 of that paper.
Each ingredient in the simulation introduces an uncertainty into the final predicted ages, luminosities, or magnitudes of the white dwarfs. The ingredients and their uncertainties are listed in Table 2, with both the original size of the uncertainties and the reduced uncertainties after fitting simultaneously the white dwarf mass distribution and local SFH in Roberts_2025. In this work, all uncertainties use the reduced sizes. Some of these uncertainties are systematic (population age, initial mass function, main sequence lifetimes, spectral evolution, merger delays, age vs kinematic relation, cooling ages, crystallisation delays) and some affect individual stars in each simulation (metallicity, and any Gaia data errors associated with each star). We refer the reader to Roberts_2025 for further details.
@@
This work focusses on two of the three methods used in Roberts_2025, comparisons of the white dwarf luminosity function and absolute Gaia G magnitude distribution, with the third - the direct age method - explored only in Appendix A for those who wish to see all three methods tested on the 100 pc samples.
@@ 3.2 Luminosity function
To calculate the bolometric magnitudes of white dwarfs in the sample, the effective temperature and radius of the white dwarf are required. This itself requires knowing the atmospheric composition of the white dwarf, i.e. hydrogen- or helium-dominated, as fitting temperature and radius requires fitting a model atmosphere and mass-radius relation to the photometry and astrometry (or, although not used in this case, fitting the same models to spectroscopy). With the 40 pc sample, all of this composition information is known, but for the 100 pc samples used in this work, this is only known for the samples of Vincent_2024 and Kilic_2025_100pc. Because of this, we chose to use temperatures and masses as derived from pure-hydrogen atmospheres for samples where no spectroscopic or incomplete spectroscopic data is known. Fig. 2 demonstrates that this assumption makes little difference, as the derived 100 pc white dwarf luminosity functions are very similar to the one for the volume complete 40 pc sample, which has full atmospheric composition information. We note that most cool white dwarfs ($T_{\rm eff}<5000\,$K) are of DC spectral type with unconstrained compositions (e.g. Elms_2022), hence the assumption of pure-hydrogen atmospheres is actually done for all samples at the faint end of the luminosity function, regardless of the availability of spectroscopy.
On the other hand, in the simulated luminosity functions, all samples without composition information rely on the 25 per cent helium fraction as described in Table 2 and each white dwarf is randomly assigned a composition. For the samples of Vincent_2024 and Kilic_2025_100pc, we can predict a slightly more accurate luminosity function using the known compositions of each white dwarf. We find that these two options make very little difference in the predicted luminosity functions.
Figure 4: Top: Results from the luminosity function method with population synthesis using all four assumed star formation forms (constant; Roberts_2025 direct age method; Fantin_2019; Mor_2019) compared to observed 100 pc white dwarf sample of GentileFusillo_2021. Error bars are the standard deviation of 100 runs of the simulations, and Poisson errors for the observed sample. Bottom: Results from the absolute Gaia G magnitude method with all four assumed star formation forms compared to the observed 100 pc sample of GentileFusillo_2021. Error bars follow the same convention as above.
@@ 3.3 Absolute Gaia G magnitude distribution
All observed magnitudes are dereddened based on either the $V$-band extinction ($A_{V}$) for that object in GentileFusillo_2021, or if missing, an $A_{V}$ extinction value sampled from a Gaussian distribution with the mean as 0.02 mag (and standard deviation of 0.015 mag), which is the average extinction over the whole 100 pc sample (GentileFusillo_2021; Sahu_2024).
Table 2: Literature and reduced uncertainties derived in Roberts_2025 as input of single-star population synthesis. The reduced uncertainties were used for all simulations in this work.
@@
Population age $\mu=10.6$ Gyr, $\sigma=0.7$ Gyr Cukanovaite_2023 $\sigma=0.5$ Gyr
Initial mass function $\mu=2.35$, $\sigma=0.1$ Salpeter_1955 $\sigma=0.075$
Initial metallicity $\mu=0.0134$, $\sigma=0.0104$ Asplund_2009 no change
Main sequence lifetimes multiplicative, $\mu=1$, $\sigma=0.048$ Byrne_2024 $\sigma=0.035$
He-atmosphere WD fraction $\mu=0.25$, $\sigma=0.1$ OBrien_2024 $\sigma=0.065$
Merger delays off/on, equal weighting Roberts_2025 off/on, 43/57 weighting
Age vs. kinematic relation multiplicative to slope of scale height, $\mu=1$, $\sigma=0.3$ Roberts_2025 no change
Cooling models multiplicative, $\mu=1$, $\sigma=0.06$ Bedard_2020 $\sigma=0.045$
Crystallisation cooling delay off/on, equal weighting Roberts_2025 no change